OpenStudy (anonymous):

What is the mass in grams of 6.5 moles of water? What equation do I use?

OpenStudy (anonymous):

for water, 1 mol of water molecules weigh (2+2+16)=18grams Important mole conversion factors: 1 mole = 6.02 x 1023 particles = formula mass in grams so 6.5 moles weigh.......??do it urself :)

OpenStudy (abb0t):

That is WRONG @AnthonyStark

OpenStudy (anonymous):

??

OpenStudy (anonymous):

oh sorry...

OpenStudy (anonymous):

hydrogen (1)....so molar mass of h2o=(1+1+16)=18

OpenStudy (abb0t):

If you're given 6.5 moles of water, all you need to do is multiply by it's molar mass of water. 6.022 \(\times\) 10\(^{23}\) is NOT the formula mass. That is avogadro's constant. If you're unfamiliar with what that is look at your book.

OpenStudy (anonymous):

sorry about that

OpenStudy (abb0t):

And yes, to fina molar mass, you add up the sum of the elements. Since you have H\(_2\), that means you have 2 \(\times\) (MASS of 1 hydrogen atom). You can find the masses on the periodic table @B.Sanders

OpenStudy (anonymous):

1 mole = 6.02 x 1023 particles = formula mass in grams sorry...by this i meant 6.02 x 1023 particles weigh formula mass in grams

OpenStudy (anonymous):

abb0t your a genius aren't you

OpenStudy (anonymous):

That's what I thought @abb0t I was just making sure. Thank you! :)

OpenStudy (abb0t):

I have NO idea what you're trying to say, @AnthonyStark but avogadro's number merely defines the number of constituent particles (atoms or molecules) in one mole of a given substance.

OpenStudy (anonymous):

i said what u just explained...in simple words..molar mass=mass of 1 mole=mass of 6.022X10^23 particles

OpenStudy (anonymous):

@B.Sanders

OpenStudy (anonymous):

this lab